OpenStudy (candy13106):

write an equation of the line that passes through the given point and is perpendicular to the given line. (5,-1), y=2x-6

OpenStudy (candy13106):

@1dyazz@dan815 @Elsa213 @piercetheveil47 @Yuba @undeadknight26 @Agent_Sniffles

OpenStudy (candy13106):

@dan815

OpenStudy (candy13106):

@Destinymasha @Data_LG2 @ShayleeCutie @gigi.antonia @Gabylovesyou @Whiterabbit541 @mathmath333 @MakaylaTracy @gold_grace123 @HaileyD @cody7744 @e_cypher @ernest34 @hhelpplzzzz @ZombieSammie

OpenStudy (anonymous):

why did u do that

OpenStudy (candy13106):

idk needed help

OpenStudy (dan815):

(5,-1), y=2x-6 \[(y+1)=\frac{-1}{2}*(x-5)\]

OpenStudy (candy13106):

uh????????????????????????????????????????????????????????????????????????????????????

OpenStudy (dan815):

\[(y+1)=\frac{-1}{2}*(x-5)\\ y=-x/2 +5/2-1\\ y=-x/2+3/2\]

OpenStudy (candy13106):

i never learned that

OpenStudy (dan815):

you have to take the egative recripocrl

OpenStudy (candy13106):

ohh now i c

OpenStudy (dan815):

negative reciprocal

OpenStudy (candy13106):

yea

OpenStudy (candy13106):

thanks

OpenStudy (anonymous):

u need help with what

OpenStudy (anonymous):

idk